COULD YOU PLEASE HELP ME SOLVE THIS EQUATION FOR Y. I have tried and I am still not getting anythere with it. Thank you.
1/2(4X -6) = 2/3(6X -9)+3

Answer by Alan3354(69443)

COULD YOU PLEASE HELP ME SOLVE THIS EQUATION FOR Y. I have tried and I am still not getting anythere with it. Thank you.
1/2(4X -6) = 2/3(6X -9)+3
2x - 3 = 4x - 6 + 3
2x - 3 = 4x - 3
2x = 4x
0 = 2x
x = 0
I don't see a Y in the equation.